AB
+BA
______
CD4
In the correctly worked addition problem above, each letter represents a different nonzero digit. What is one possible value of the two-digit number represented above as AB ?
(A) 74 (B) 31 (C) 86 (D) 47 (E) 95
option: (c)86

Found 2 solutions by richwmiller, ravi5:
Answer by richwmiller(17219)

We know the sum must end in 4
74+47 ends in 1
31+13 ends in 4 but has only two digits
86+68 ends in 4 and has three digits sum is 154
47+74 ends in 1 same as A)
95+59 ends in 4 and has three digits sum is 154
A,B, C and D are all unique so 95 and 59 can't use the 5 again in the sum
So the only acceptable answer is 86 and 68.

Answer by ravi5(2)

AB
+ BA
------
CD4
Soln:put the value of A & B
A=6 or 8,B=8 or 6 because B+A=4
therefore,8+6 or 6+8 =14 then carry 1
Again 1+A+B=D then D=5 then carry 1
and C=1
ans:68 or 86
Option:(c)86